# Ithaca, NY 14853, USA

### v1

@@ 1. Introduction
The combinatorial nonpositive curvature condition of $m$-location for simplicial complexes was introduced by Osajda in [16] as a method for proving hyperbolicity for $3$-manifolds. While local systolicity of a complex depends only on the combinatorics around a single vertex, $m$-location depends on the combinatorics around pairs of adjacent vertices, allowing more flexibility. Osajda proved that simply connected $8$-located locally $5$-large complexes are Gromov hyperbolic and used this to obtain a new solution to a problem of Thurston. The second named author studied a version of $8$-location, suggested by Osajda [16, Subsection 5.1] and showed that simply connected, $8$-located simplicial complexes are Gromov hyperbolic [13]. In later work. the second named author introduced another combinatorial curvature condition, called the $5/9$-condition, and showed that the complexes which satisfy it, are $8$-located and therefore Gromov hyperbolic as well [14].
@@
A simply connected $7$-located locally $5$-large complex has quadratic isoperimetric function. Consequently, the fundamental group of a $7$-located locally $5$-large complex has quadratic isoperimetric function
@@ 2. Preliminaries
A $(k,l)$-dwheel $W=W_{1}\cup W_{2}$ in $X$ is the union of two wheels $W_{1}=(w_{l};v_{1},v_{2},v_{3}=w_{l-1},\ldots,$ $v_{k})$ and $W_{2}=(v_{2};w_{1},w_{2},\ldots,w_{l})$ with $v_{1}=w_{1}$ or $v_{1}\sim w_{1}$. See Figure 1. If $v_{1}=w_{1}$, we call $W$ a planar dwheel. If $v_{1}\sim w_{1}$, we call $W$ a nonplanar dwheel. The boundary $\partial W$ of $W$ is the cycle $(w_{1},w_{2},\ldots,w_{l-1}=v_{3},v_{4},\ldots,v_{k})$. The boundary $\partial W$ is the subcomplex of $W$ induced on its vertices excluding the wheel centers $w_{l}$ and $v_{2}$. Note that $\partial W$ has length $k+l-4$ if $W$ is planar and $k+l-3$ if it is nonplanar.
@@
For the purposes of this paper, we need only refer to local weak systolicity, a condition which together with simple connectedness is equivalent to weak systolicity by the local-to-global theorem of Chepoi and Osajda [8, Theorem A]. In order to define local weak systolicity, we first need to introduce some terminology. An extended $5$-wheel $\widehat{W}_{5}=(c;v_{1},v_{2},v_{3},v_{4},v_{5};a)$ in a flag simplicial complex $X$ is a subcomplex consisting of a $5$-wheel $W_{5}=(c;v_{1},v_{2},v_{3},v_{4},v_{5})$ together with an additional vertex $a$ that spans a triangle $\langle a,v_{1},v_{2}\rangle$ with $v_{1}$ and $v_{2}$ but that is not adjacent to any other vertex of $W_{5}$.
@@ 3. Weakly systolic complexes are $7$-located
Consider $W_{1}\cup\{w_{2}\}$. Note that $w_{2}\sim v_{1}$ and $w_{2}\sim v_{2}$ and that, since $W_{2}$ is a full $k$-wheel, the vertices $w_{2}$ and $w_{5}$ are distinct and $w_{2}\nsim w_{5}$ and $w_{2}\nsim v_{3}$. If we had $w_{2}\sim v_{4}$ then $(w_{2},v_{2},w_{5},v_{4})$ would be a $4$-cycle and then $5$-largeness of $X$ would contradict fullness of either $W_{1}$ or $W_{2}$. So $w_{2}\nsim v_{4}$. We also have $w_{2}\nsim v_{5}$ arguing otherwise similarly using the $4$-cycle $(w_{2},v_{2},w_{5},v_{5})$. Thus $W_{1}\cup\{w_{2}\}=(w_{5};v_{1},v_{2},v_{3},v_{4},v_{5};w_{2})$ is an extended $5$-wheel. Thus, by the $\widehat{W_{5}}$-condition, there is a vertex $\bar{w}_{2}$ that is adjacent to but distinct from every vertex of $W_{1}\cup\{w_{2}\}$. If $\bar{w}_{2}=w_{3}$ then $W_{1}\cup W_{2}=W_{1}\cup\{w_{2},w_{3}\}$ is contained in the $1$-ball centered at $\bar{w}_{2}$ and we are done. So we may assume that $\bar{w}_{2}\notin W_{1}\cup W_{2}$. By a similar argument, there is a vertex $\bar{w}_{3}\notin W_{1}\cup W_{2}$ that is adjacent to every vertex of $W_{1}\cup\{w_{3}\}$. If $\bar{w}_{2}=\bar{w}_{3}$ then $W_{1}\cup W_{2}=W_{1}\cup\{w_{2},w_{3}\}$ is contained in the $1$-ball centered at $\bar{w}_{2}$ and we are done. So we may assume that $\bar{w}_{2}\neq\bar{w}_{3}$. Then $(v_{1},\bar{w}_{2},v_{3},\bar{w}_{3})$ is a $4$-cycle and $5$-largeness of $X$ and fullness of $W_{1}$ imply that $\bar{w}_{2}\sim\bar{w}_{3}$. Finally, the $5$-largeness of $X$ in consideration of the $4$-cycle $(\bar{w}_{2},w_{2},w_{3},\bar{w}_{3})$ implies that $W_{1}\cup W_{2}$ is contained in the $1$-ball centered at either $\bar{w}_{2}$ or $\bar{w}_{3}$. ∎
@@ 4. A $\operatorname{CAT}(0)$ metric for $7$-located disc diagrams
Since $D$ is simply connected, it suffices to show that the sum of the angles of the corners of triangles incident to any interior vertex $v$ of $D$ is at least $2\pi$ [4, Theorem II.5.4 and Lemma II.5.6]. Because $D$ is flag, it has no interior vertices of degree $3$. Any interior vertex of degree $4$ or $5$ is the central vertex of a flattened wheel and so has angle sum exactly $2\pi$. Away from flattened wheel centers, corner angles of triangles are either $\frac{\pi}{4}$, $\frac{3\pi}{10}$ or $\frac{\pi}{3}$. Thus any interior vertex of degree at least $8$ has angle sum at least $2\pi$.
Thus we may assume that $v$ is an interior vertex of degree $6$ or $7$. If $v$ is not incident to any flattened wheel then its incident triangles are all regular Euclidean and so have corner angle $\frac{\pi}{3}$ so that $v$ has angle sum $\frac{6\pi}{3}=2\pi$ or $\frac{7\pi}{3}>2\pi$. So we may assume that $v$ is incident to a flattened wheel. Since $\deg(v)\leq 7$, by Lemma 4.1, any central vertex of a flattened wheel $W$ incident to $v$ has degree $5$. Then $v$ is incident to a vertex of degree $5$ and any triangle corner incident to $v$ has angle at least $\frac{3\pi}{10}$. So, by Lemma 4.1, we have $\deg(v)=7$ and so $v$ has angle sum at least $\frac{21\pi}{10}>2\pi$. ∎
@@
The area of a minimal disc diagram for a loop of length $n$ in a $\operatorname{CAT}(0)$ space is bounded by the area of a circle of circumference $n$ in the Euclidean plane [4, Theorem III.2.17]. It follows that if $D$ is a $7$-located disc of boundary length $n$, it has area at most $\frac{n^{2}}{4\pi}$ when endowed with the flattened wheel metric. Each triangle of $D$ has area at least $\frac{1}{4}$ so the number of triangles of $D$ is at most $f(n)=\frac{n^{2}}{\pi}$. ∎

### v2

@@ 1. Introduction
The combinatorial nonpositive curvature condition of $m$-location for simplicial complexes was introduced by Osajda in [14] as a method for proving hyperbolicity for $3$-manifolds. While local systolicity of a complex depends only on the combinatorics around a single vertex, $m$-location depends on the combinatorics around pairs of adjacent vertices, allowing more flexibility. Osajda proved that simply connected $8$-located locally $5$-large complexes are Gromov hyperbolic and used this to obtain a new solution to a problem of Thurston. The second named author studied a version of $8$-location, suggested by Osajda [14, Subsection 5.1] and showed that simply connected, $8$-located simplicial complexes are Gromov hyperbolic [11]. In later work, the second named author introduced another combinatorial curvature condition, called the $5/9$-condition, and showed that the complexes which satisfy it, are $8$-located and therefore Gromov hyperbolic as well [12].
@@
A simply connected $7$-located locally $5$-large complex has quadratic isoperimetric function. Consequently, the fundamental group of a $7$-located locally $5$-large complex has quadratic isoperimetric function.
@@ 2. Preliminaries
A $(k,l)$-dwheel $W=W_{1}\cup W_{2}$ in $X$ is the union of two wheels $W_{1}=(w_{l};v_{1},v_{2},v_{3},\ldots,$ $v_{k})$ and $W_{2}=(v_{2};w_{1},w_{2},\ldots,w_{l})$ with $v_{3}=w_{l-1}$ and either $v_{1}=w_{1}$ or $v_{1}\sim w_{1}$. See Figure 1. If $v_{1}=w_{1}$, we call $W$ a planar dwheel. If $v_{1}\sim w_{1}$, we call $W$ a nonplanar dwheel. The boundary $\partial W$ of $W$ is the cycle $(w_{1},w_{2},\ldots,w_{l-1}=v_{3},v_{4},\ldots,v_{k})$. The boundary $\partial W$ is the subcomplex of $W$ induced on its vertices excluding the wheel centers $w_{l}$ and $v_{2}$. Note that $\partial W$ has length $k+l-4$ if $W$ is planar and $k+l-3$ if it is nonplanar.
@@
$a$$c$$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$
Figure 2. An extended $5$-wheel.
For the purposes of this paper, we need only refer to local weak systolicity, a condition which together with simple connectedness is equivalent to weak systolicity by the local-to-global theorem of Chepoi and Osajda [6, Theorem A]. In order to define local weak systolicity, we first need to introduce some terminology. An extended $5$-wheel $\widehat{W}_{5}=(c;v_{1},v_{2},v_{3},v_{4},v_{5};a)$ in a flag simplicial complex $X$ is a subcomplex consisting of a $5$-wheel $W_{5}=(c;v_{1},v_{2},v_{3},v_{4},v_{5})$ together with an additional vertex $a$ that spans a triangle $\langle a,v_{1},v_{2}\rangle$ with $v_{1}$ and $v_{2}$ but that is not adjacent to any other vertex of $W_{5}$. See Figure 2.
@@ 3. Weakly systolic complexes are $7$-located
Consider $W_{1}\cup\{w_{2}\}$. Note that $w_{2}\sim v_{1}$ and $w_{2}\sim v_{2}$. Also, since $W_{2}$ is a full $k$-wheel, the vertices $w_{2}$ and $w_{5}$ are distinct and $w_{2}\nsim w_{5}$ and $w_{2}\nsim v_{3}$. If we had $w_{2}\sim v_{4}$ then $(w_{2},v_{2},w_{5},v_{4})$ would be a $4$-cycle and then $5$-largeness of $X$ would contradict fullness of either $W_{1}$ or $W_{2}$. So $w_{2}\nsim v_{4}$. We also have $w_{2}\nsim v_{5}$ arguing otherwise similarly using the $4$-cycle $(w_{2},v_{2},w_{5},v_{5})$. Thus $W_{1}\cup\{w_{2}\}=(w_{5};v_{1},v_{2},v_{3},v_{4},v_{5};w_{2})$ is an extended $5$-wheel. Thus, by the $\widehat{W_{5}}$-condition, there is a vertex $\bar{w}_{2}$ that is adjacent to but distinct from every vertex of $W_{1}\cup\{w_{2}\}$. If $\bar{w}_{2}=w_{3}$ then $W_{1}\cup W_{2}=W_{1}\cup\{w_{2},w_{3}\}$ is contained in the $1$-ball centered at $\bar{w}_{2}$ and we are done. So we may assume that $\bar{w}_{2}\notin W_{1}\cup W_{2}$. By a similar argument, there is a vertex $\bar{w}_{3}\notin W_{1}\cup W_{2}$ that is adjacent to every vertex of $W_{1}\cup\{w_{3}\}$. If $\bar{w}_{2}=\bar{w}_{3}$ then $W_{1}\cup W_{2}=W_{1}\cup\{w_{2},w_{3}\}$ is contained in the $1$-ball centered at $\bar{w}_{2}$ and we are done. So we may assume that $\bar{w}_{2}\neq\bar{w}_{3}$. Then $(v_{1},\bar{w}_{2},v_{3},\bar{w}_{3})$ is a $4$-cycle and $5$-largeness of $X$ and fullness of $W_{1}$ imply that $\bar{w}_{2}\sim\bar{w}_{3}$. Finally, the $5$-largeness of $X$ in consideration of the $4$-cycle $(\bar{w}_{2},w_{2},w_{3},\bar{w}_{3})$ implies that $W_{1}\cup W_{2}$ is contained in the $1$-ball centered at either $\bar{w}_{2}$ or $\bar{w}_{3}$. ∎
@@ 4. A $\operatorname{CAT}(0)$ metric for $7$-located disc diagrams
Since $D$ is simply connected, it suffices to show that the sum of the angles of the corners of triangles incident to any interior vertex $v$ of $D$ is at least $2\pi$ [3, Theorem II.5.4 and Lemma II.5.6]. Because $D$ is flag, it has no interior vertices of degree $3$. Any interior vertex of degree $4$ or $5$ is the central vertex of a flattened wheel and so has angle sum exactly $2\pi$. Away from flattened wheel centers, corner angles of triangles are either $\frac{\pi}{4}$, $\frac{3\pi}{10}$ or $\frac{\pi}{3}$. Thus any interior vertex of degree at least $8$ has angle sum at least $2\pi$.
Thus we may assume that $v$ is an interior vertex of degree $6$ or $7$. If $v$ is not incident to any flattened wheel then its incident triangles are all regular Euclidean and so have corner angle $\frac{\pi}{3}$. Thus $v$ has angle sum $\frac{6\pi}{3}=2\pi$ or $\frac{7\pi}{3}>2\pi$. So we may assume that $v$ is incident to a flattened wheel. Since $\deg(v)\leq 7$, by Lemma 4.1, any central vertex of a flattened wheel $W$ incident to $v$ has degree $5$. Then $v$ is incident to a vertex of degree $5$ and any triangle corner incident to $v$ has angle at least $\frac{3\pi}{10}$. So, by Lemma 4.1, we have $\deg(v)=7$ and so $v$ has angle sum at least $\frac{21\pi}{10}>2\pi$. ∎
@@
The area of a minimal disc diagram for a loop of length $n$ in a $\operatorname{CAT}(0)$ space is bounded by the area of a circle of circumference $n$ in the Euclidean plane [3, Theorem III.2.17]. It follows that if $D$ is a $7$-located disc of boundary length $n$, it has area at most $\frac{n^{2}}{4\pi}$ when endowed with the flattened wheel metric. Each triangle of $D$ has area at least $\frac{1}{4}$ so the number of triangles of $D$ is at most $f(n)=\frac{n^{2}}{\pi}$. ∎